Two large metal plates of area 0.7m2 face each other. They are 5.2 cm apart and have equal but opposite chargeson their inner surfaces. If the magnitude E of theelectric field between the plates is 53 N/C, what is the magnitude of the charge oneach plate? Neglect edge effects.
Answer #1
Given a parallel plate capacitor set up
Plate area A = 0.7 m2
El field between th plates, E = 53 N/C
Find charge Q on each plate.
We know the el field is related to surface charge density,σ = Q/A
In fact, E = σ/εo and so, σ =εo .E
Therefore, charge,
           Q= σ.A =  εo .E.A
              = 8.854x10-12 x 53 x 0.7
              = 3.28x10-10 C =328 pC